OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today?

OpenStudy (anonymous):

wow this one again! where oh where does this come from? get at least one a week.

OpenStudy (anonymous):

\[x+y=840\]\[3x-2y=870\] does that look right?

OpenStudy (callisto):

If I interpret the question correctly ..... Let x = number of one-step equations y = number of two-step equations ''The number of one-step equations and two-step equations who have been eliminated today is equal to 840!'' => x+y = 840 -(1) ''three times the number of one-step equations minus twice the number of two-step equations is equal to 870'' => 3x - 2y = 870 - (2) (1)x2 + (2) 2(x+y) + (3x - 2y) = 2(840) + 870 All you need to do is to solve x.

OpenStudy (anonymous):

yeah looks good, but i still want to know where this question comes from. does everyone in the world take the same on line class? and what the monkey is a "one step" equation? and "math idol judges"? puhleez

OpenStudy (anonymous):

I got is from my online class. So would that answer be 330, 510, 520, or 660?

OpenStudy (callisto):

It's better for you to solve it...

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

solve \[x+y=840\] for \(y\) get \[y=840-x\] substitute in to second equation with careful use of parentheses get \[3x-2(840-x)=870\] and finally solve that linear equation for \(x\) write what you think the answer is